OpenStudy (anonymous):

Write the slope-intercept form of the equation of the line that passes through the points (1,-5) and (4,-1).

OpenStudy (bahrom7893):

M=(-1+5)/(4-1)=4/3

OpenStudy (anonymous):

find slope: m = (-1-(-5))/(4-1) = 4/3 y-(-5) = (4/3)(x-1) y+5 = (4/3)x-(4/3) y = (4/3)x +11/3